Multiply
(c+1/3)(c+1/2)
Thank you

Answer by Alan3354(69443)

Multiply
(c+1/3)(c+1/2)
----------------
= c*(c+1/2) + 1/3(c+1/2)
= c^2 + c/2 + c/3 + 1/6
= c^2 + 5c/6 + 1/6
Multiply by each term, then collect terms. That's all.